# Why Can’t I Open My Drawer? Mitigating Object-Driven Shortcuts in Zero-Shot Compositional Action Recognition

### v1

@@ B.2 Implementation details on RCORE
Table 6: Hyperparameters used for training RCORE on each dataset.
Config Sth-com EK100-com
Optimizer AdamW [20]
Base learning rate (visual) 5e-4 1e-4
Base learning rate (text) 1e-4 1e-5
@@
$\alpha,\beta,\gamma,\delta$ 0.2, 1.0, 1.0, 0.8
$p_{aug}$ 0.5 0.3
@@
(start, end) epoch for scaling $\delta$ (15, 20)
Scale factor for $\lambda$ in VOCAMix 0.2 0.5
K for top-k selection in Margin loss 10
@@
Training epochs 30
Table 7: Top-10 failure cases of the existing ZS-CAR model on unseen compositions of the Sth-com test set. We employ C2C [16] trained on Sth-com [16] as baseline for analysis. We use ‘sth’ instead of ‘something’ to shorten. We report the results with unbiased open-world inference scores. ‘Prediction Rate’ denotes the proportion of mispredictions classified as the specific composition. ‘Verb/Object-side’ indicates the most frequently co-occurring counterpart for each mispredicted component in the training set, along with its co-occurrence ratio. The top-10 cases listed below account for 9.5% of the total mispredictions on unseen compositions.
@@
Table 8: Top-10 failure cases of the existing ZS-CAR model on unseen compositions of the EK100-com test set. We employ C2C [16] trained on EK100-com as baseline for analysis. We report the results with unbiased open-world inference scores. ‘Prediction Rate’ denotes the proportion of mispredictions classified as the specific composition. ‘Verb/Object-side’ indicates the most frequently co-occurring counterpart for each mispredicted component in the training set, along with its co-occurrence ratio. The top-10 cases listed below account for $\sim$ 39% of the total mispredictions on unseen compositions.
@@
Appendix C Additional Evidence of Object-Driven Shortcuts
@@
In Table 7 and Table 8, we investigate top-10 failure cases of the model trained with the existing ZS-CAR learning approach [16]. The top-10 failure cases shown in the Table 7 constitute 9.5% of the misclassifications on unseen compositions for Sth-com [16], compared to 49% for EK100-com as shown in Table 8. For the analysis of Sth-com [16], we exclude ambiguous cases arising from annotation issues to obtain meaningful results. A prime example is the top-1 failure case, (‘Covering [something] with [something]’, ‘Paper’), where 66% of the errors are misclassified as ‘Pencil’. Upon visual inspection, we confirm that nearly all these videos actually depict the action of covering a pencil with paper. In Table 7 and Table 8, we observe a common tendency across both datasets: the model frequently misclassifies verbs as their temporally reversed or opposing counterparts. Specifically, we demonstrate that these mispredictions stem from the model’s reliance on co-occurrence statistics. For instance, the model misclassifies unseen ‘(Closing, Drawer)’ samples as ‘(Opening, Drawer)’, which accounts for 36% of the errors. Notably, ‘Opening’ is the verb most frequently paired with ‘Drawer’ in the training set. These findings suggest that the model has failed to learn robust verb representations that generalize to unseen compositions, largely due to object-driven shortcut learning.
@@
Table 9: Additional ablation study of RCORE on the Sth-com. We provide the additional ablation study results on Sth-com [16]. In every experiment, we use C2C [16] with CLIP-B/16 [28] as our baseline. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
(a) Effects of TORC with different temporal aggregation.
(b) Effects top-$K$ selection for Margin loss.
@@
Seen Unseen Seen Unseen Seen ($\Delta_{\mathrm{CG}}$) Unseen ($\Delta_{\mathrm{CG}}$) H.M.
@@
$K$ Verb Object Composition FCP
Seen Unseen Seen Unseen Seen Unseen H.M.
10 64.60 54.42 66.99 55.68 46.71 29.80 36.38 8.81
100 63.63 53.93 67.58 55.66 46.43 29.18 35.84 9.89
3451 62.75 53.63 66.89 55.20 45.32 29.16 35.49 8.02
(c) Effects of hyperparameters of VOCAMix.
(d) Effects of hyperparameters of TORC.
$p_{aug}$ Verb Object Composition
Seen Unseen Seen Unseen Seen Unseen H.M.
0.0 63.60 54.36 67.72 56.10 46.31 30.08 36.47
0.3 64.67 54.83 68.18 55.37 47.77 29.93 36.80
0.5 64.26 54.65 67.32 56.61 46.86 30.35 36.84
1.0 64.02 53.80 67.04 55.26 46.30 29.07 35.72
@@
$\gamma$ scaling scaling Verb Object Composition
start epoch end epoch Seen Unseen Seen Unseen Seen Unseen H.M.
0.0 0 0 63.60 54.36 67.72 56.10 46.31 30.08 36.47
0.5 1 1 65.59 55.63 68.49 54.80 48.33 30.19 37.17
1.0 1 1 64.88 55.81 67.21 55.12 47.06 30.69 37.15
1.0 5 5 64.93 56.03 67.67 55.81 47.20 31.38 37.70
1.0 5 10 65.65 56.80 68.59 55.19 48.65 31.35 38.13
(e) Effects of hyperparameters of Margin loss.
@@
$\delta$ scaling scaling Verb Object Composition
start epoch end epoch Seen Unseen Seen Unseen Seen Unseen H.M.
0.3 10 10 63.09 52.94 67.90 54.86 46.31 28.48 35.27
0.3 15 15 63.96 53.70 67.70 55.68 46.57 29.28 35.96
0.3 15 20 64.01 53.76 67.53 55.61 46.87 29.42 36.15
0.5 15 20 63.85 53.61 67.90 56.32 46.79 29.49 36.18
0.8 15 20 64.60 54.42 66.99 55.68 46.71 29.80 36.38
Appendix D Additional Results
D.1 Comprehensive ablation results
In Table 9, we provide the comprehensive ablation study results to examine each design choice of RCORE, using the Sth-com [16] dataset.
Effects of TORC loss with different temporal aggregation methods.
In Table 9 (d), we demonstrate that the weak temporal order modeling caused by object-driven shortcut learning cannot be resolved merely by employing more sophisticated temporal aggregation methods. We compare the baseline’s temporal average pooling with attention pooling (Attn. pool). In attention pooling, a linear layer computes frame-wise attention weights from the input verb feature sequences, aggregating them via a weighted sum to obtain the final verb feature for classification. While attention pooling yields a slight improvement in verb accuracy, it fails to reduce the compositional gap ($\Delta_{CG}$) on unseen compositions. In contrast, our TORC loss achieves significant improvements in H.M. with both aggregation methods. Notably, the combination of temporal average pooling and TORC loss enables the model to overcome the negative compositional gap.
Effects of top-$K$ selection for Margin loss.
In Table 9 (e), we investigate the impact of the hyperparameter $K$, which determines the number of top-ranked logits to which the Margin loss is applied. We find that applying the margin selectively to the frequent-but-incorrect logits within the top-10 ($K=10$) reports the highest H.M., outperforming the approach of applying it to all seen composition logits ($K=3451$) (36.4% vs. 35.5%). This suggests that penalizing only the most confusing candidates—specifically those that frequently co-occur with the input components—effectively mitigates co-occurrence bias without compromising the model’s overall performance. For instance, given an input like ‘(Pretending to tear, Paper)’, our intention is to penalize plausible but incorrect distractors such as ‘(Tearing, Paper)’ or ‘(Folding, Paper)’, while avoiding penalties on irrelevant classes like ‘(Throwing, Paper)’ that are already ranked low.

### v2

@@ 0.C.2 Implementation details on RCORE
Table 7: Hyperparameters used for training RCORE on each dataset. We denote InternVideo2 [wang2024iv2] as ‘IV2’.
Config Sth-com [li2024c2c] EK100-com
CLIP [radford2021clip] IV2 [wang2024iv2] CLIP [radford2021clip] IV2 [wang2024iv2]
Optimizer AdamW [adamw]
Base learning rate (visual) 5e-4 5e-5 1e-4 5e-5
Base learning rate (text) 1e-4 5e-5 1e-5 5e-5
@@
$\alpha,\beta,\gamma,\delta$ 0.2, 1.0, 1.0, 0.1
$p_{\text{CPR}}$ 0.95 0.8 0.95 0.8
@@
(start, end) epoch for scaling $\delta$ (15, 15) (30, 35) (35, 40) (30, 35)
Scale factor for $\lambda$ in CPR 0.5
@@
Training epochs 40 50 40 40
Experimental Setup. We implement our method using PyTorch and conduct all experiments on 8 NVIDIA Tesla V100 GPUs. For video preprocessing, we uniformly sample frames from each video to generate clips of size $3\times 16\times 224\times 224$ (channels $\times$ frames $\times$ height $\times$ width) for the CLIP [radford2021clip] backbone and $3\times 8\times 224\times 224$ for the InternVideo2 [wang2024iv2] backbone.
Model Architecture. We utilize CLIP-B/16 [radford2021clip] as our visual/text backbone. Following prior work [li2024c2c], we adopt AIM [yang2023aim], a parameter-efficient spatio-temporal adapter for CLIP [radford2021clip], as the video encoder. We also employ InternVideo2-CLIP-B/14 [wang2024iv2] as our modern video-pretrained VLM backbone and utilize LoRA [hu2022lora] for parameter-efficient fine-tuning. We use CoOp-style [zhou2022coop] learnable text prompts for all models. We also construct the verb encoder with two temporal convolution layers with ReLU activations, while the object encoder employs temporal average pooling followed by a two-layer MLP with ReLU activations.
To extract $T$ frame-level features with the InternVideo2 [wang2024iv2] backbone, we modify its attention pooling mechanism. Originally, the model concatenates a CLS token with video tokens in $\mathbb{R}^{NT\times D}$, where $N$ is the number of spatial patches and $D$ is the feature dimension, and applies attention pooling to produce a single video-level representation in $\mathbb{R}^{1\times D}$. In our approach, we adapt this to operate at the frame level. Specifically, for each frame, we concatenate the CLS token with the corresponding spatial tokens in $\mathbb{R}^{N\times D}$. This creates $T$ separate feature sets, each in $\mathbb{R}^{(N+1)\times D}$. Finally, we apply attention pooling independently to each of these sets, resulting in the frame-level video features in $\mathbb{R}^{T\times D}$.
Training Strategy. We train all models, including the baselines and RCORE, following the hyperparameters detailed in Table 7, using a total batch size of 128. For CPR, we employ the unsupervised foreground estimation algorithm FAME [ding2022fame] to identify high-motion regions. FAME [ding2022fame] isolates the moving foreground from background regions based on frame differences and color statistics. To generate the mixed samples, we first sample the mixing coefficient from a $\text{Beta}(2.0,2.0)$ distribution, following prior work [yun2019cutmix, zhang2018mixup]. We then scale it by 0.5 to obtain the final blending coefficient $\lambda$. We apply this CPR’s augmentation to the samples within a batch with a probability of $p_{\text{aug}}$. For the frequent verb-object pairs to apply the $L_{\text{CPR}}$ loss term, we select pairs that possess a strong co-occurrence prior from both verb and object perspectives as the frequent pairs $\mathbb{Y}_{\text{freq}}$, which is defined as $\mathbb{Y}_{\text{freq}}=\mathbb{Y}_{\text{freq}}^{O\mid V}\cap\mathbb{Y}_{\text{freq}}^{V\mid O}$. $\mathbb{Y}_{\text{freq}}^{O\mid V}$ is the set of pairs where the conditional probability of an object given a verb in the training data is greater than $\mu+\sigma$. Here, $\mu$ and $\sigma$ represent the mean and standard deviation of the conditional probabilities calculated on the training set. The $\mathbb{Y}_{\text{freq}}^{V\mid O}$ is defined similarly for the conditional probability of a verb given an object. $(\mu_{O|V}+\sigma_{O|V},\mu_{V|O}+\sigma_{V|O})$ are calculated as $(0.13,0.19)$ for the Sth-com [li2024c2c] dataset and $(0.23,0.40)$ for EK100-com.
Loss configuration. For the total loss, we set the weights for the verb and object components to $\alpha=0.2$ and for the composition loss to $\beta=1.0$, following prior work [li2024c2c]. As demonstrated in Table 11 (d), we empirically find it beneficial to linearly warm up the weights for the TORC loss: $\gamma$ is increased from 0.0 to 1.0 over 5 epoch to 10 epoch. We use $\delta$, the loss scale of $L_{\text{CPR}}$, differently depending on the dataset and backbone, as detailed in Table 7.
Inference. During inference, we adopt a single-view, single-crop protocol. To select the best model during training and for bias calibration, we use the harmonic mean of the accuracies for seen and unseen compositions as the model selection criterion.
Table 8: Top-10 failure cases of the existing ZS-CAR model on unseen compositions of the Sth-com test set. We employ C2C [li2024c2c] trained on Sth-com [li2024c2c] as baseline for analysis. We use ‘sth’ instead of ‘something’ to shorten. We report the results with unbiased open-world inference scores. ‘Prediction Rate’ denotes the proportion of mispredictions classified as the specific composition. ‘Verb/Object-side’ indicates the most frequently co-occurring counterpart for each mispredicted component in the training set, along with its co-occurrence ratio. The top-10 cases listed below account for 9.5% of the total mispredictions on unseen compositions.
@@
Table 9: Top-10 failure cases of the existing ZS-CAR model on unseen compositions of the EK100-com test set. We employ C2C [li2024c2c] trained on EK100-com as baseline for analysis. We report the results with unbiased open-world inference scores. ‘Prediction Rate’ denotes the proportion of mispredictions classified as the specific composition. ‘Verb/Object-side’ indicates the most frequently co-occurring counterpart for each mispredicted component in the training set, along with its co-occurrence ratio. The top-10 cases listed below account for $\sim$ 39% of the total mispredictions on unseen compositions.
Rank Ground Truth (Verb, Object) Error Top-2 Misprediction (seen/unseen) Prediction Most Frequent Co-occurring Component
@@
Appendix 0.D Additional evidence of object-driven shortcuts
@@
Failure case analysis of the existing ZS-CAR model. In Table 8 and Table 9, we investigate top-10 failure cases of the model trained with the state-of-the-art method [li2024c2c]. The top-10 failure cases shown in the Table 8 constitute 9.5% of the misclassifications on unseen compositions for Sth-com [li2024c2c], compared to 49% for EK100-com as shown in Table 9. For the analysis of Sth-com [li2024c2c], we exclude ambiguous cases arising from annotation issues to obtain meaningful results. A prime example is the top-1 failure case, (‘Covering [something] with [something]’, ‘Paper’), where 66% of the errors are misclassified as ‘Pencil’. Upon visual inspection, we confirm that nearly all these videos actually depict the action of covering a pencil with paper. In Table 8 and Table 9, we observe a common tendency across both datasets: the model frequently misclassifies verbs as their temporally reversed or opposing counterparts. Specifically, we demonstrate that these mispredictions stem from the model’s reliance on co-occurrence statistics. For instance, the model misclassifies unseen ‘(Closing, Drawer)’ samples as ‘(Opening, Drawer)’, which accounts for 36% of the errors. Notably, ‘Opening’ is the verb most frequently paired with ‘Drawer’ in the training set. These findings suggest that the model has failed to learn robust verb representations that generalize to unseen compositions, largely due to object-driven shortcut learning.
Figure 13: Similarity analysis between original and reversed verb features. We plot the cosine similarity between the original and reversed verb features for the baseline (C2C [li2024c2c]) and RCORE with the CLIP [radford2021clip] backbone on the Sth-com [li2024c2c] dataset. The baseline maintains a high similarity (+0.92) during entire training, revealing limited temporal sensitivity. In contrast, the similarity becomes strongly negative for RCORE as training progresses, indicating improved temporal discriminative capability.
Similarity analysis of verb features. In Figure 13, we track the cosine similarity between the original verb features $\mathbf{f}^{V}$ and the reversed verb features $\mathbf{f}^{V}_{\text{rev}}$ for the baseline [li2024c2c] over the course of training. The baseline [li2024c2c] maintains a high cosine similarity (+0.92+0.92), revealing its reliance on object-driven shortcuts rather than temporally grounded verb representations. Additionally, we plot the similarity of RCORE in Figure 13. RCORE drives this similarity to a strongly negative value (0.79-0.79), indicating that the model successfully distinguishes opposite temporal semantics, e.g. ‘opening’ vs. ‘closing’.
Appendix 0.E Additional Results
0.E.1 Macro Compositional Gap
Table 10: Macro Compositional Gap results. We show Macro Compositional Gap ($\Delta_{\text{CG}}^{\text{macro}}$) results of C2C [li2024c2c] and RCORE on the Sth-com [li2024c2c] and EK100-com datasets. The best numbers are highlighted.
Backbone Method Sth-com EK100-com
Seen Unseen Seen Unseen
CLIP AIM [yang2023aim] +1.18 -0.47 +0.95 -1.10
C2C [li2024c2c] +0.78 +0.16 +1.56 -0.38
RCORE +0.53 +0.64 -0.03 +0.11
InternVideo2 LoRA [hu2022lora] 0.00 -0.16 +0.60 -0.53
C2C [li2024c2c] +0.59 +0.14 +0.83 -0.38
RCORE -0.33 +0.76 +0.02 +0.85
In this section, we define Macro Compositional Gap ($\Delta_{\text{CG}}^{\text{macro}}$). $\Delta_{\text{CG}}^{\text{macro}}$ is the macro version of Compositional Gap ($\Delta_{\text{CG}}$), which averages $\Delta_{\text{CG}}$ per composition class, ensuring robustness to class imbalance. Let $\mathbb{Y}_{\text{test}}^{C}\subseteq\mathbb{Y}^{V}\times\mathbb{Y}^{O}$ denote the set of unique composition labels that appear in the test split. For each $\mathbf{y}^{C}=(\mathbf{y}^{V},\mathbf{y}^{O})\in\mathbb{Y}_{\text{test}}^{C}$, let $\mathrm{Acc}^{C}_{\mathbf{y}^{C}}$ be the top-1 composition accuracy restricted to samples of $\mathbf{y}^{C}$, and let $\mathrm{Acc}^{V}_{\mathbf{y}^{C}}$ and $\mathrm{Acc}^{O}_{\mathbf{y}^{C}}$ be the corresponding verb/object accuracies on the same samples of $\mathbf{y}^{C}$, computed from the joint top-1 composition predictions. Then we define
$\Delta_{\text{CG}}^{\text{macro}}=\frac{1}{|\mathbb{Y}^{C}_{\text{test}}|}\sum_{\mathbf{y}^{C}\in\mathbb{Y}^{C}_{\text{test}}}\left[\text{Acc}^{C}_{\mathbf{y}^{C}}-\left(\text{Acc}^{V}_{\mathbf{y}^{C}}\times\text{Acc}^{O}_{\mathbf{y}^{C}}\right)\right].$ (14)
In Table 10, we report $\Delta_{\text{CG}}^{\text{macro}}$ performances of baselines and RCORE on the Sth-com [li2024c2c] and EK100-com datasets. RCORE consistently achieves the highest $\Delta_{\text{CG}}^{\text{macro}}$ on unseen compositions across different backbones and datasets, demonstrating its robust compositional reasoning capabilities.
0.E.2 Additional ablation results
In Table 11, we provide the comprehensive ablation study results to examine each design choice of RCORE, using the Sth-com [li2024c2c] dataset.
Effects of TORC with different temporal aggregation methods. In Table 11 (a), we demonstrate that the weak temporal order modeling caused by object-driven shortcut learning cannot be resolved merely by employing more sophisticated temporal aggregation methods. We compare the baseline’s temporal average pooling with attention pooling (Attn. pool). In attention pooling, a linear layer computes frame-wise attention weights from the input verb feature sequences, aggregating them via a weighted sum to obtain the final verb feature for classification. While attention pooling yields a slight improvement in verb accuracy, it fails to reduce the compositional gap ($\Delta_{CG}$) on unseen compositions. In contrast, our TORC achieves significant improvements in harmonic mean (H.M.) with both aggregation methods. Notably, the combination of temporal average pooling and TORC enables the model to overcome the negative compositional gap.
@@ 0.E.2 Additional ablation results
Table 11: Additional ablation study of RCORE on Sth-com. We provide the additional ablation study results on Sth-com [li2024c2c]. In every experiment, we use C2C [li2024c2c] with CLIP-B/16 [radford2021clip] as our baseline. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.). The best numbers are highlighted.
(a) Effects of TORC with different
temporal aggregation.
(b) Effects of input samples in CPR.
@@
V@S V@U O@S O@U Seen ($\Delta_{\mathrm{CG}}$) Unseen ($\Delta_{\mathrm{CG}}$) H.M.
@@
Inputs Verb Object Composition
V@S V@U O@S O@U Seen Unseen H.M.
Original 62.54 52.97 56.90 57.05 37.43 31.41 34.16
Mixup [zhang2018mixup] 62.37 52.56 63.50 59.20 41.62 31.45 35.83
CPR w/o F.E. [ding2022fame] 63.23 54.29 65.15 56.49 43.57 30.86 36.13
CPR w/ F.E. [ding2022fame] 64.17 54.55 63.60 58.70 42.74 33.05 37.28
(c) Effects of hyperparameters of CPR. $p_{\text{CPR}}$ Verb Object Composition V@S V@U O@S O@U Seen Unseen H.M. 0.8 64.66 52.89 65.06 58.77 45.20 31.14 36.88 0.9 64.70 55.50 64.73 57.69 44.75 32.60 37.72 0.95 63.95 54.05 63.97 59.52 43.09 33.22 37.51 1.0 64.95 55.50 61.49 57.20 41.82 32.84 36.79
@@
(d) Effects of hyperparameters of TORC. $\gamma$ scaling scaling Verb Object Composition start epoch end epoch V@S V@U O@S O@U Seen Unseen H.M. 0.5 1 1 65.59 55.63 68.49 54.80 48.33 30.19 37.17 1.0 1 1 64.88 55.81 67.21 55.12 47.06 30.69 37.15 1.0 5 5 64.93 56.03 67.67 55.81 47.20 31.38 37.70 1.0 5 10 65.65 56.80 68.59 55.19 48.65 31.35 38.13
@@
(e) Effects of hyperparameters of $L_{\text{CPR}}$. $m$ $\delta$ scaling scaling Verb Object Composition start epoch end epoch V@S V@U O@S O@U Seen Unseen H.M. 0.5 0.1 15 15 67.06 58.13 63.67 58.66 45.00 35.16 39.48 0.5 0.1 15 20 66.16 58.78 63.75 57.96 44.65 34.89 39.17 0.5 0.5 15 15 66.16 57.03 62.75 57.56 43.55 33.77 38.04 0.5 0.5 15 20 67.04 56.77 61.74 57.56 43.37 34.30 38.31 1.0 0.1 15 15 66.31 58.56 63.77 57.99 44.47 34.64 38.94 1.0 0.1 15 20 66.88 58.24 63.88 57.01 44.90 34.21 38.83
Effects of input samples in CPR. In Table 11 (b), we validate the effects of input sample types in CPR. We argue that conventional image-based augmentations are insufficient for representing newly created unseen composition labels. For example, Mixup [zhang2018mixup] blends verbs and objects concurrently, failing to preserve the temporal difference information that is essential for verb recognition. As a result, these methods yield a lower verb@unseen-comp than the baseline (52.6% vs. 53.0%). In contrast, our CPR adopts a static frame mixing strategy, aiming to generate novel compositions by mixing only object information while preserving the verbs, shows modest performance gains over the baseline for both verb and object accuracy on unseen compositions, resulting in the best H.M. Furthermore, we observe that mixing a static frame into the estimated foreground region shows additional performance boost (+0.2 points in the H.M.) over applying it to the entire background, as it creates more realistic samples. We denote the latter as ‘CPR w/o F.E.’, indicating ‘without foreground estimation’.
Table 12: Ablation study on EK100-com. We provide the results of the ablation study on EK100-com presented in the main paper. In every experiment, we use C2C [li2024c2c] with CLIP-B/16 [radford2021clip] as our baseline. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
(a) Effects of CPR and TORC. CPR TORC Verb Object Composition) V@S V@U O@S O@U Seen Unseen H.M. 66.19 49.71 56.61 47.48 42.72 22.38 29.38 66.06 51.28 54.23 50.56 38.93 27.10 31.96 65.90 52.31 57.28 46.74 43.48 23.67 30.65 67.52 54.08 54.00 49.04 40.13 27.40 32.56
(b) Effects of each TORC loss term. $L_{\cos}$ $L_{\text{ent}}$ Verb Object Composition V@S V@U O@S O@U Seen Unseen H.M. 66.19 49.71 56.61 47.48 42.72 22.38 29.38 65.43 52.11 56.54 47.44 42.67 24.24 30.92 66.00 51.06 56.64 47.37 42.82 22.53 29.53 65.90 52.31 57.28 46.74 43.48 23.67 30.65
(c) Effects of $L_{\text{CPR}}$ in CPR. $L_{\text{CPR}}$ Verb Object Composition V@S V@U O@S O@U Seen Unseen H.M. 66.35 50.79 54.72 50.16 39.25 26.60 31.71 66.06 51.28 54.23 50.56 38.93 27.10 31.96
Effects of hyperparameters of RCORE. In Table 11 (c), we show the effect of $p_{\text{CPR}}$, the per-batch probability for CPR application, revealing a trade-off between seen and unseen composition accuracies. Decreasing $p_{\text{CPR}}$ slightly from 1.0 (e.g. 0.95 or 0.9) yields optimal unseen accuracy improvements with minimal drops in seen performance, whereas lowering it further (e.g. $p_{\text{CPR}}=0.8$) degrades the gains on unseen pairs. Consequently, we tune $p_{\text{CPR}}$ for each dataset and backbone based on the best validation H.M. of seen and unseen accuracies, as shown in Table 7. In Table 11 (d), we observe that introducing the explicit temporal order modeling task via TORC starting from 5 epoch, after initial training, is effective. In particular, the strategy of linearly increasing the loss weight from epoch 5 to 10 is the most effective. Finally, in Table 11 (e), we analyze the impact of the loss weights of $L_{\text{CPR}}$ in CPR. Increasing the margin $m$ or loss weight $\delta$ degrades performance, suggesting that excessive training penalties are detrimental. We find $m=0.5$ and $\delta=0.1$ to be optimal. Contrary to the results in Table 11 (d), maintaining a constant low loss weight ($\delta=0.1$) for $L_{\text{CPR}}$ yields better performance than applying progressive weight scaling.
0.E.3 Ablation studies on EK100-com
In Table 12, we provide the ablation study results conducted on the EK100-com dataset, complementing the results discussed with the Sth-com [li2024c2c] dataset in the main paper. In Table 12 (a), we demonstrate that while CPR and TORC are individually effective, integrating both components yields the best performance on EK100-com, outperforming the baseline in H.M. by 3.2 points. In Table 12 (b), using only $L_{cos}$ in TORC yields the highest H.M. on EK100-com. In Table 12 (c), surpressing co-occurrence priors with $L_{cos}$ improves H.M. by 0.3 points, consistent with the results on Sth-com [li2024c2c].
0.E.4 Results with validation-set-tuned bias calibration
Table 13: Sth-com results with bias calibration. We show the top-1 verb, object, and composition classification accuracies (%) on the Sth-com [li2024c2c] test set. We tune the bias term for unseen composition predictions on the validation set and apply the selected bias (‘Best Bias’) to the test set. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.). The best numbers are highlighted.
[…]
In this section, we provide the open-world biased results with validation-set-tuned calibration. Bias calibration is a standard technique in compositional zero-shot learning [purushwalkam2019taskdriven, mancini2021compcos, kim2023hierarchical, nayak2023csp, wu2025cond, li2024context] and ZS-CAR [li2024c2c, jiang2025dhd, jung2025crr, ye2025logic]. It compensates for the inherent issue where logits for unseen compositions—not optimized during training—are naturally lower than those for seen compositions. Specifically, it adds a scalar bias term exclusively to the unseen composition logits. Previous ZS-CAR works [li2024c2c, jiang2025dhd, jung2025crr, ye2025logic] often tuned this bias using test-set ground truth under a closed-world setting to report the ‘best’ seen and unseen accuracies. We argue that this approach is highly unrealistic.
Instead, we adopt an open-world validation-set-tuned bias calibration. We tune the bias on the validation set using open-world logits and then apply it to the test set. This approach remains applicable to real-world scenarios while effectively improving unseen composition performance as a post-processing step.
In Table 13, we present the calibrated results of the baselines and RCORE on the Sth-com [li2024c2c] dataset. Notably, the calibrated performance of RCORE is identical to its original unbiased performance. This indicates that our training label space expansion strategy via our proposed CPR is effective. Unlike baseline methods that overfit solely to seen compositions, it successfully provides meaningful supervision for unseen compositions during training. Furthermore, the unbiased performance of RCORE consistently outperforms the best calibrated performance of all baselines. This demonstrates that RCORE is highly suitable for real-world applications, delivering strong ZS-CAR performance without the need for any post-processing.
Table 14: Sth-com results with InternVideo2-1B. We show the top-1 verb, object, and composition classification accuracies (%) on the Sth-com [li2024c2c] and EK100-com datasets with the InternVideo2-1B [wang2024iv2] backbone. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.). The best numbers are highlighted.
Sth-com results
Method Verb Object Composition
@Seen Comp @Unseen Comp H.M. @Seen Comp @Unseen Comp H.M. Seen ($\Delta_{\text{CG}}$) Unseen ($\Delta_{\text{CG}}$) H.M.
C2C [li2024c2c] 67.50 58.21 62.51 70.58 61.62 65.80 50.23 (+2.59) 34.96 (-0.91) 41.23
RCORE 67.49 60.47 63.79 68.83 63.63 66.13 48.27 (+1.82) 38.35 (-0.13) 42.74
EK100-com results
Method Verb Object Composition
@Seen Comp @Unseen Comp H.M. @Seen Comp @Unseen Comp H.M. Seen ($\Delta_{\text{CG}}$) Unseen ($\Delta_{\text{CG}}$) H.M.
C2C [li2024c2c] 66.95 52.22 58.67 60.57 50.81 55.26 45.91 (+5.36) 26.66 (+0.13) 33.73
RCORE 66.36 52.88 58.86 56.83 52.10 54.36 41.86 (+4.16) 28.79 (+1.24) 34.12